Can we exchange the denominator of lhs with numerator of rhs

Grade:10

1 Answers

satish
7 years ago
yes of course for exaample we have an expression   2x+3/4-5x=5/9  so we can transpose denominator of lhs to the numerator of rhs by 2x+3/5=4-5x/9